OpenStudy (anonymous):

HELP TO EARN A MEDAL PLEASE !!! Create a table for the exponential function f(x)=3*2^x. PLEASE HELP !!!!!

OpenStudy (jdoe0001):

pick a few random "x values" and get the "y" from that, put them in a table

OpenStudy (zale101):

note: f(x) is y so if you put a number for x, solve the equation to get f(x) which is y keep doing that until you have a set of pairs to point them on the graph

OpenStudy (jdoe0001):

\(\large{ f(x)=3\cdot 2^{\color{red}{ x}}\qquad \qquad \begin{array}{llll} x&y\\ \hline\\ 0&3\cdot 2^{\color{red}{ 0}}\implies 3\cdot 1\implies 3\\ 1&3\cdot 2^{\color{red}{ 1}}\implies 3\cdot 2\implies 6\\ ...&... \end{array}}\)

OpenStudy (anonymous):

@jdoe0001 okay what if i use the ''x values '' as -2 , -1,0,1,2 they how would my table turn out?

OpenStudy (zale101):